OpenStudy (anonymous):

solve each equation. check your solution: 10/3x + 4/3 = 7+x/2x

OpenStudy (saifoo.khan):

\[\frac{10}{3} x + \frac 43 = \frac{7 + x}{2x}\]This?

OpenStudy (anonymous):

yeah

OpenStudy (saifoo.khan):

The "x" is in numerator or denominator?

OpenStudy (anonymous):

ok so the answer is 7+x/2x?

OpenStudy (anonymous):

@alexwee is that continued?

OpenStudy (alexwee123):

yep and thats the answer

OpenStudy (anonymous):

ok wait whats the answer?

OpenStudy (alexwee123):

x=21/41

OpenStudy (anonymous):

ok thanks

OpenStudy (alexwee123):

oh wait i might have made a mistake

OpenStudy (anonymous):

ok

OpenStudy (alexwee123):